OpenStudy (anonymous):

You go with your little brother to a drag strip where cars race. He is into this stuff and brings his stopwatch. The purple car starting from rest travels a quarter of a mile (1320.0 ft) in 8.06 seconds. What was the acceleration? What would the equation be?

OpenStudy (anonymous):

d= Vi*t +1/2 *a *t^2 if Vi =0 : d= 1/2 *a *t^2 1320 = 1/2*a*(8.06)^2

OpenStudy (fellowroot):

You go with your little brother to a strip bar.....and he is into his stuff.....what?